# More Expressive Attention with Negative Weights

### v2

@@ Abstract
We propose a novel attention mechanism, named Cog Attention, that enables attention weights to be negative for enhanced expressiveness, which stems from two key factors: (1) Cog Attention can shift the token deletion and copying function from a static OV matrix to dynamic QK inner products, with the OV matrix now focusing more on refinement or modification. The attention head can simultaneously delete, copy, or retain tokens by assigning them negative, positive, or minimal attention weights, respectively. As a result, a single attention head becomes more flexible and expressive. (2) Cog Attention improves the model’s robustness against representational collapse, which can occur when earlier tokens are over-squashed into later positions, leading to homogeneous representations. Negative weights reduce effective information paths from earlier to later tokens, helping to mitigate this issue. We develop Transformer-like models which use Cog Attention as attention modules, including decoder-only models for language modeling and U-ViT diffusion models for image generation. Experiments show that models using Cog Attention exhibit superior performance compared to those employing traditional softmax attention modules. Our approach suggests a promising research direction for rethinking and breaking the entrenched constraints of traditional softmax attention, such as the requirement for non-negative weights. Our code is available at https://github.com/trestad/CogAttn.
Attention with Negative Weights
@@ 1 Introduction
The Transformer architecture (Vaswani et al., 2017) has achieved success across numerous applications, such as language modeling (Brown et al., 2020) and image generation (Dosovitskiy et al., 2021). A crucial factor contributing to its success is the softmax attention mechanism (Bahdanau et al., 2015).
@@
Figure 1: In the Indirect Object Identification (IOI) task (Wang et al., 2023), a language model should identify the indirect object ($IO$) from a context that includes both $IO$ and a subject ($S$). Subfigures (a) and (b) show how Cog Attention and softmax attention process a single sequence through a process of elimination, but in different ways: a softmax attention head with a deletion-function OV matrix eliminates all attended tokens. While the $IO$ token receives less attention than $S$, it is also deleted. In contrast, Cog Attention shifts functions like deletion or copying from a static OV matrix to dynamic query-key inner products, allowing the head to assign negative weights to $S$ tokens for elimination while preserving $IO$s. Subfigures (c) and (d) show the attention of these two heads on names across the entire dataset vs the direction of their output embeddings. Cog Attention exhibits a weaker correlation between the attention weights assigned to $IO$s and the extent of their deletion. For further details, please see Section 3.
@@
Softmax ensures non-negative attention weights, but we argue that it limits the expressiveness of the attention mechanism. Figure 1 shows one possible way that negative attention weights enhance the model’s expressiveness using the same number of parameters: in a softmax attention head, the QK matrix (Elhage et al., 2021) determines the relevant tokens for attention, while the OV matrix governs the processing of these attended tokens (e.g., deletion or copying). Suppose a softmax attention head has an OV matrix capable of deleting tokens that the QK matrix attends to; since attention weights must be non-positive, a useful token is also somewhat deleted. By allowing negative attention weights, however, deletion or copying can be expressed through the sign of the attention weight and accomplished during the weighted summation of value vectors. This functional shift also allows the OV matrix to focus more on higher-level tasks, such as refinement or modification, rather than solely handling contextual deletions or copies. Consequently, negative attention weights eliminate irrelevant tokens while preserving useful ones, mitigating the risk of “friendly fire” on useful tokens.
@@
Despite the potential benefits of incorporating negative weights in attention mechanisms, this question has been rarely explored. Apart from the common belief that attention weights should naturally be non-negative, introducing negative weights can lead to challenges such as training instability, numerical overflow, and difficulties in attention normalization due to issues like division by zero, etc.
@@
In this paper, we propose a novel attention mechanism named Cog Attention [^1] The name is derived from the attention pattern, which resembles cogs. See Figure 10. that enables negative weights. Cog Attention exhibits superior properties from a mechanistic interpretability perspective and surpasses softmax attention in various applications without introducing any additional parameters. In Section 3, we provide mechanistic evidence for the expressiveness of Cog Attention:
@@
(1) We identify attention heads that share the same working mechanism as exemplified above (Figures 1(b) and (d)), which shift the contextual process from the static OV matrix to dynamic QK inner products, with the OV matrix focusing more on refinement or modification. Irrelevant tokens are assigned negative weights for elimination, while other tokens are well preserved at the same time. This demonstrates Cog Attention’s enhanced flexibility and expressiveness compared to softmax attention.
@@
(2) We demonstrate that models using Cog Attention exhibit improved robustness against representational collapse (Liu et al., 2020; Xie et al., 2023). Representational collapse refers to the phenomenon where representations become homogeneous in the later positions of a sequence within deep Transformer models. Barbero et al. (2024) contended that this issue arises because earlier tokens are “over-squashed” into later positions as the layer goes deeper. The negative weights in Cog Attention reduce the effective information paths from earlier tokens to later positions, thereby alleviating over-squashing and, consequently, mitigating representational collapse.
@@
In Section 4, we develop Transformer-like models that use Cog Attention as attention modules and evaluate their performance across various tasks. Specifically, we train decoder-only language models for language modeling, and U-ViT diffusion models (Bao et al., 2023) for both unconditional and text-conditioned image generation. Our results show that across a wide range of tasks, these models equipped with Cog Attention achieve improved performance over the vanilla Transformer architecture using softmax attention.
@@ 2 Method
We begin by presenting the formulation of our proposed Cog Attention. Then, we discuss the design motivation and underlying principles.
@@ 2.1 Formulation
Let $\textbf{q},\textbf{k},\textbf{v}\in\mathbbm{R}^{n\times d}$ represent the query, key, and value vectors in an attention head. $n$ is the number of input tokens and $d$ is the hidden states dimension. The general attention computation can be expressed as follows:
@@
where $\textbf{p}_{i}\in\mathbbm{R}^{1\times n}$ is the $i$-th row of the inner-product matrix. $\textbf{a}_{i}\in\mathbbm{R}^{1\times n}$ is the $i$-th row of the attention weights. $\textbf{o}_{i}\in\mathbbm{R}^{1\times d}$ is the weighted summation of vectors attended. [^2] Eq.1 is a causal attention formulation, as a token $i$ can only attend to preceding tokens $j\leq i$.
$\phi(\cdot)$ is the softmax function in a traditional attention module:
@@
[…]
The subtraction of $m_{i}$ in the index aims to avoid numerical overflow.
@@
$\phi(\textbf{p}_{i})_{j}$ $=\frac{\texttt{SignExp}(\textbf{p}_{i,j})}{\sum^{i}_{k=0}\lvert\texttt{SignExp}(\textbf{p}_{i,k})\rvert},$ (3)
$\text{where}\ \texttt{SignExp}(\textbf{p}_{i,j})$ $=s_{i,j}\cdot\texttt{exp}(s_{i,j}\cdot\textbf{p}_{i,j}-m_{i}),$
$m_{i}=\texttt{max($\lvert\textbf{p}_{i}\rvert$)}\ \ \text{and}$ $\ \ s_{i,j}=\texttt{sign}(\textbf{p}_{i,j}).$
This formulation enables Cog Attention to produce negative attention weights. In the following subsection, we introduce and explain several design motivations and the underlying principles.
@@ 2.2 Design Principle
(1) The way to introduce negative weights. Although the inner product of query and key vectors naturally contains both positive and negative values, we apply an exponential function to this inner product and subsequently recover the sign of each term. This method is driven by our observation that an effective attention pattern for convergence must demonstrate sufficient kurtosis; i.e., it should be sparse and sharp enough. Without the exponential function, the attention pattern tends to be too flat, which can impede training convergence. We also tried using a cubic function as an alternative, which would eliminate the sign recovery process while still offering attention weights with adequate kurtosis. However, we ultimately chose the exponential function because of its convenience in gradient computation.
(2) The way to avoid numerical overflow. In Eq.3, we avoid numerical overflow in exponential functions by subtracting the maximum absolute value from the index. This approach differs from the softmax, which subtracts the maximum value, as seen in Eq.2. Our approach ensures that the maximum input to $\texttt{SignExp}(\cdot)$ remains 0, effectively avoiding overflow caused by a large $s_{i,j}\cdot\textbf{p}_{i,j}$, as illustrated in Figure 2.
Figure 2: The subtraction of the maximum absolute value in a row of query-key inner products avoids numerical overflow.
@@
# q, k, and v’s shape: n x d
# mask shape: n x n, 0 for tokens to be masked, 1 for others
@@
def Cog_Attention_fast(q, k, v, mask):
@@
Figure 3: A naive implementation of Cog Attention in Pytorch, alongside an equivalent yet faster implementation.
(3) The way to normalization. In softmax functions, the denominator in Eq.2 is the summation of the numerator, serving to normalize the outputs. This process ensures that the resulting attention weights sum to 1, with each term constrained within the range of $[0,1]$. Previous studies suggested that, for better convergence, the sum of attention weights in a row should remain stable, although not necessarily equal to a constant 1 (Wortsman et al., 2023). If not, training tricks are necessary to remain convergence and training stability (Ramapuram et al., 2024).
However, we challenge these findings. As shown in Eq. 3, normalizing by summing all outcomes of $\texttt{SignExp}(\cdot)$ may result in a zero denominator, causing NaN errors. To address this, we propose using the sum of the absolute values of the outcomes of $\texttt{SignExp}(\cdot)$ as the denominator. This adjustment leads to a non-constant summation across a row in the attention weight matrices. Nevertheless, our experiments demonstrate that Cog Attention maintains training stability and does not hinder convergence. This is because: (1) our method ensures that $\sum_{j=0}^{i}\lvert\phi(\textbf{p}_{i})_{j}\rvert$ remains constant at 1; and (2) based on the observation that the expectation of v is close to zero, adding or subtracting these value vectors—assumed to follow a multivariate Gaussian distribution—does not disrupt the norm expectation of the results, i.e., $\textbf{o}_{i}$ in Eq. 1.
We present a naive PyTorch implementation of Cog Attention, as shown in Figure 3, which closely follows the above description. Furthermore, we offer a more efficient equivalent implementation.
@@ 3 Enhanced Expressiveness of Cog Attention: A Mechanistic Interpretability Perspective
In this section, we provide mechanistic interpretability evidence to demonstrate that negative attention weights can enhance neural networks’ expressiveness.
3.1 Cog Attention Enhances the Flexibility of Attention Heads
Due to its unconstrained attention weights, Cog Attention enables processes such as deletion or copying, transitioning from a static OV matrix to dynamic query-key products. This capability stands as a key advantage of Cog Attention, facilitating concurrent processes within a single head, thereby enhancing the models’ flexibility and expressiveness.
We trained a Transformer language model with 141 million parameters and a Transformer-like language model using Cog Attention of the same size, respectively. Details regarding the model training can be found in Section 4. We studied attention heads’ working mechanisms on the indirect object identification (IOI) task (Wang et al., 2023), where the model is given with a context that includes the names of two people. For instance, given the input “Christopher and David had a lot of fun at school. David gave a ring to,” “Christopher” is the indirect object ($IO$), while “David” is the subject ($S$). The correct answer in IOI task is always the $IO$, which in this case is “Christopher.” There are 100 samples in the dataset.
To identify the most influential attention heads contributing to correct predictions in each model, we employed the path patching algorithm (Wang et al., 2023). We identified two significant heads: the 4th Cog Attention head in Layer 9 ($\text{CH}_{9.4}$) and the 11th softmax head in Layer 9 ($\text{SH}_{9.11}$). These heads accomplish the IOI task through a process of elimination. Figure 1(a) and (b) illustrate the attention weight patterns for these heads. Additionally, we computed the attention weights from the final token to both the $IO$ and $S$ tokens, plotted against the inner product of the heads’ outputs and $IO$ and $S$’s individual embedding, as shown in Figure 1(c) and (d).
@@ 3.1 Cog Attention Enhances the Flexibility of Attention Heads
In contrast, $\text{CH}_{9.4}$ assigns negative weights to $S$, effectively eliminating it while assigning minimal weights to $IO$s. Even if an $IO$ is assigned with a negative weight, we contend that the OV matrix in $\text{CH}_{9.4}$ is relieved from deletion to perform post-processing such as refinement or modification, and thus retains $IO$s. Evidence for this claim comes from the eigenvalue positivity [^3] Defined in the “Summarizing OV/QK Matrices” section in  (Elhage et al., 2021) for the OV matrices in two heads. The eigenvalue positivity, which indicates the OV matrix’s tendency toward copying (close to 1), deletion (close to -1), or abstract post-processing such as refinement or modification (close to 0), is 0.78 for $\text{CH}_{9.4}$ and -0.95 for $\text{SH}_{9.11}$. Additionally, the low Pearson correlation coefficient of $0.13$ between the attention weights and the inner product suggests little to no correlation between these two variables. As a result, $IO$s are not under “friendly fire” from $\text{CH}_{9.4}$.
3.2 Cog Attention is More Robust to Representational Collapse
Representational collapse (Liu et al., 2020; Xie et al., 2023) refers to that the representations at many positions become similar as the Transformer model goes deeper. Barbero et al. (2024) proposed an explanation for this issue, suggesting that earlier input tokens have more information paths to the final position compared to later tokens. As a result, the information from earlier tokens “over-squashed” into the representation at the final position, which is used for next-token prediction. This over-squashing causes representational collapse, leading Transformer models to struggle with distinguishing contexts that differ only slightly.
@@ 3.2 Cog Attention is More Robust to Representational Collapse
Figure 5: The $L_{+\infty}$ norms are computed for Transformer language models utilizing Cog Attention (subfigures (a) and (b)) and softmax attention subfigures (c) and (d)), respectively. The results are normalized by dividing them by the values obtained when $n=200$. Cog Attention makes language models more robust to representational collapse.
We demonstrate that Transformer models using Cog Attention are more robust to representational collapse than those using softmax attention. This improvement may be attributed to the negative weights reducing the number of effective information paths from earlier tokens to later positions, thereby mitigating the degree of over-squashing.
To assess the extent of representational collapse in a Transformer-based language model, we evaluate two tasks. In Task 1, “Finding a Zero,” the model processes two input sequences: the first consists of $n+1$ ones, and the second consists of $n$ ones with a leading zero. Following Barbero et al. (2024), we compute the $L_{+\infty}$ norm of the differences between the representations at the final positions of the two sequences. A larger $L_{+\infty}$ norm indicates that the model more effectively distinguishes between the two contexts, reflecting less representational collapse. Task 2, “Counting Ones,” follows a similar evaluation process.
We evaluate our trained 141M models using Cog Attention and softmax attention, respectively. The value of $n$ is set to 200, 400, 600, 800, 1000, and 2000, respectively.
Figure 5 illustrates the results. For each value of $n$ and each task, models employing Cog Attention exhibit significantly higher relative $L_{\infty}$ norms compared to those using softmax attention, highlighting Cog Attention’s robustness to representational collapse. This advantage of Cog Attention could be beneficial for tasks that involve long and complex contexts, such as retrieval-augmented generation (RAG) (Lewis et al., 2020).
Model ARC-E ARC-C PIQA SST-2 MNLI MRPC QQP RTE Avg.
Transformer 42.34 19.54 57.73 51.72 33.21 68.63 36.82 51.99 45.24
Cogformer 43.90 19.54 59.09 54.59 34.12 68.38 37.00 52.71 46.16
Table 1: Cogformer achieves a higher average accuracy on various tasks compared with Transformer models.

### v3

@@ Abstract
We propose a novel attention mechanism, named Cog Attention, that enables attention weights to be negative for enhanced expressiveness. This stems from two key factors: (1) Cog Attention enhances parameter flexibility. For example, unlike traditional softmax attention heads, which use a static output-value (OV) matrix to delete or copy inputs that the heads attend to, Cog Attention naturally learns to use the sign of dynamic query-key (QK) inner products to represent these operations. This enables Cog Attention to perform multiple operations simultaneously within a single head. Meanwhile, Cog Attention’s OV matrix can focus more on refinement. (2) Cog Attention improves the model’s robustness against representational collapse by preventing earlier tokens from “over-squashing” into later positions. We develop Transformer-like models that use Cog Attention as attention modules, including decoder-only models with up to 1 billion parameters for language modeling and U-ViT diffusion models for image generation. Experiments show that models using Cog Attention exhibit superior performance compared to those employing traditional softmax attention modules. Our approach suggests a promising research direction for rethinking and breaking the entrenched constraints of traditional softmax attention, such as the requirement for non-negative weights.
Attention mechanism, Language models, Negative attention weights
@@ 1 Introduction
The Transformer architecture (Vaswani et al., 2017) has achieved success across numerous applications, such as language modeling (Brown et al., 2020) and image generation (Dosovitskiy et al., 2021). A key factor in its success is the softmax attention mechanism (Bahdanau et al., 2015).
@@
Figure 1: In the Indirect Object Identification (IOI) task (Wang et al., 2023), a language model should identify the indirect object ($IO$) from a context that includes both the $IO$ and a subject ($S$). Figures (a) and (b) illustrate how Cog Attention and softmax attention perform IOI through a process of elimination: a softmax attention head with a deletion-function OV matrix eliminates all attended tokens. While the $IO$ token receives less attention than $S$, it is also deleted. In contrast, Cog Attention shifts functions like deletion or copying from a static OV matrix to dynamic query-key inner products, allowing the head to assign negative weights to $S$ tokens for elimination while preserving the $IO$s. Figures (c) and (d) show attention weights on names versus the direction of the heads’ output across the entire dataset. Cog Attention preserves the $IO$s better. For further details, please see Section 3.
@@
Softmax ensures non-negative attention weights, but we argue that it limits the expressiveness of the attention mechanism. Figure 1 shows one way in which negative attention weights can enhance the model’s expressiveness while using the same number of parameters: in a softmax attention head, the query-key (QK) matrix (Elhage et al., 2021) determines the relevant tokens for attention, while the output-value (OV) matrix governs the processing of these attended tokens (e.g., deletion or copying). Suppose a softmax attention head has an OV matrix capable of deleting tokens attended to by the QK matrix; since attention weights must be non-positive, a useful token would also be somewhat deleted. By allowing negative attention weights, however, deletion or copying can be expressed through the sign of the attention weight and accomplished during the weighted summation of value vectors. This functional shift also allows the OV matrix to focus more on higher-level tasks, such as refinement or modification, rather than solely handling contextual deletions or copies. Consequently, the risk of “friendly fire” on useful tokens is mitigated.
@@
Despite the potential benefits of incorporating negative weights in attention mechanisms, this question has been rarely explored. Apart from the common belief that attention weights should naturally be non-negative, introducing negative weights can lead to challenges such as training instability, numerical overflow, and difficulties in attention normalization due to issues like division by zero.
@@
In this paper, we propose a novel attention mechanism named Cog Attention [^1] The name is derived from the attention pattern, which resembles cogs (see Figure 11). that enables negative weights. Cog Attention exhibits superior properties in terms of its working mechanisms and outperforms softmax attention in various applications, without introducing additional parameters or hyperparameters. In Section 3, we provide mechanistic interpretation of its enhanced expressiveness:
@@
(1) We identify several Cog Attention heads in our pre-trained language models that share the same working mechanism as exemplified above (Figures 1(b) and (d)), which shift the contextual process from the static OV matrix to dynamic QK inner products, with the OV matrix focusing more on refinement or modification. Irrelevant tokens are assigned negative weights for elimination, while other tokens are preserved. This demonstrates Cog Attention’s enhanced flexibility and expressiveness compared to softmax attention.
@@
(2) We empirically and theoretically demonstrate that models using Cog Attention exhibit improved robustness against representational collapse (Liu et al., 2020; Xie et al., 2023). Representational collapse refers to the phenomenon where representations become homogeneous, especially in the later positions of a sequence, within deep Transformer models. Barbero et al. (2024) contended that this issue arises because earlier tokens are “over-squashed” into later positions as the layer depth increases. The negative weights in Cog Attention reduce the effective information paths from earlier tokens to later positions, thereby alleviating over-squashing and, consequently, mitigating representational collapse.
@@
In Section 4, we develop Transformer-like models that use Cog Attention as attention modules. Specifically, we train decoder-only language models with parameter scales ranging from 140M to 1B for language modeling, and U-ViT diffusion models (Bao et al., 2023) for both unconditional and text-conditioned image generation. Our results show that, across a wide range of tasks, these models equipped with Cog Attention achieve improved performance compared to the vanilla Transformer architecture using softmax attention. Additionally, in Section 5, we discuss some important properties of Cog Attention that could be beneficial for future work.
@@ 2 Method
We begin by presenting the formulation of our proposed Cog Attention, followed by a discussion of the design motivation and underlying principles, as well as how to efficiently implement it.
@@ 2.1 Formulation
Let $\textbf{q},\textbf{k},\textbf{v}\in\mathbbm{R}^{T\times d}$ represent the query, key, and value vectors in an attention head. $T$ is the number of input tokens, and $d$ is the dimension of the hidden states. The general attention computation can be expressed as follows:
@@
where $\textbf{p}_{i}\in\mathbbm{R}^{1\times T}$ is the $i$-th row of the inner-product matrix. $\textbf{a}_{i}\in\mathbbm{R}^{1\times T}$ is the $i$-th row of the attention weights. $\textbf{o}_{i}\in\mathbbm{R}^{1\times d}$ is the weighted sum of attended vectors. [^2] Eq.1 is a causal attention formulation, as a token $i$ can only attend to preceding tokens $j\leq i$.
$\phi(\cdot)$ is softmax function in a traditional attention module:
@@
where $m_{i}=\texttt{max($\textbf{p}_{i}$)}$. The subtraction of $m_{i}$ in the exponent aims to avoid numerical overflow.
@@
$\phi(\textbf{p}_{i})_{j}=\frac{\texttt{SignExp}(\textbf{p}_{i,j})}{\sum^{i}_{k=0}\lvert\texttt{SignExp}(\textbf{p}_{i,k})\rvert},\text{ where}$ (3)
$\texttt{SignExp}(\textbf{p}_{i,j})=s_{i,j}\cdot\texttt{exp}(s_{i,j}\cdot\textbf{p}_{i,j}-\tilde{m}_{i}),$
$\tilde{m}_{i}=\texttt{max($\lvert\textbf{p}_{i}\rvert$)}\ \ \text{\normalsize{and}}\ \ s_{i,j}=\texttt{sign}(\textbf{p}_{i,j}).$
This formulation enables negative attention weights. In the following subsection, we introduce the design motivations and explain the underlying principles.
@@ 2.2 Design Principle
(1) The way to introduce negative weights. Although the inner product of query and key vectors naturally contains both positive and negative values, we apply an exponential function to this inner product and subsequently recover the sign of each term. This method is driven by our observation that an effective attention pattern for convergence must demonstrate sufficient kurtosis—that is, it should be sparse and sharp enough. Without the exponential function, the attention pattern tends to be too flat, which can impede training convergence. We also tried using a cubic function as an alternative, which would eliminate the sign recovery process while still offering attention weights with adequate kurtosis. However, we ultimately chose the exponential function because of its convenience in gradient computation.
(2) The way to avoid numerical overflow. In Eq.3, we avoid numerical overflow in exponential functions by subtracting the maximum absolute value from the exponent. This differs from softmax, which subtracts the maximum value, as seen in Eq.2. Our approach ensures that the maximum input to $\texttt{SignExp}(\cdot)$ remains 0, effectively avoiding overflow caused by a large $s_{i,j}\cdot\textbf{p}_{i,j}$, as shown in Figure 2.
Additionally, subtracting the maximum absolute value effectively preserves the relative magnitude between negative and positive inner products in the final attention weights. For instance, as shown in Figure 2(b), if the minimum inner product is –100 and the maximum is 1, we expect the final attention weights to reflect that the absolute value of the smallest negative weight still exceeds that of the highest positive weight. Conversely, Figure 2(a) illustrates the opposite scenario, where positive weights dominate.
Figure 2: The subtraction of the maximum absolute value from a row of query-key inner products avoids numerical overflow. Meanwhile, this approach maintains the relative importance of negative and positive inner products in the final attention weights.
@@
# q, k, and v’s shape: T x d
# mask shape: T x T, 0 for tokens to be masked, 1 for others
@@
def Cog_Attention_faster(q, k, v, mask):
@@
Figure 3: A naive implementation of Cog Attention in Pytorch, alongside an equivalent yet faster implementation. By writing a fused kernel in Triton (Tillet et al., 2019), Cog Attention achieves the same efficiency as softmax attention.
(3) The way to normalization. In the softmax function, the denominator in Eq. 2 is the sum of the numerators, serving to normalize the outputs. This process ensures that the resulting attention weights sum to 1, with each term constrained within the range of $[0,1]$. Previous studies suggested that, for better convergence, the sum of attention weights in a row should remain stable, although not necessarily equal to a constant 1 (Wortsman et al., 2023). If not, training tricks are necessary to maintain convergence and training stability (Ramapuram et al., 2024).
However, we challenge these beliefs. As shown in Eq. 3, normalizing by summing all outcomes of $\texttt{SignExp}(\cdot)$ may result in a zero denominator, causing NaN errors. To address this, we propose using the sum of the absolute values of the outcomes of $\texttt{SignExp}(\cdot)$ as the denominator. This adjustment leads to a non-constant summation across a row in the attention weight matrices. Nevertheless, our experiments demonstrate that Cog Attention maintains training stability and does not hinder convergence. This is because: (1) $\sum_{j=0}^{i}\lvert\phi(\textbf{p}_{i})_{j}\rvert$ remains constant at 1; and (2) based on the observation that the expectation of v during pre-training is close to zero, adding or subtracting these value vectors—assumed to follow a multivariate Gaussian distribution—does not disrupt the norm expectation of the results, i.e., $\textbf{o}_{i}$ in Eq. 1.
2.3 Efficiency
Figure 3 presents a naive PyTorch implementation of Cog Attention, which closely follows the above description. Furthermore, we offer a more efficient equivalent version in PyTorch. Compared to softmax attention, this version still requires additional sign and absolute value operations. In practice, when implementing Cog Attention with a fused kernel in Triton (Tillet et al., 2019), the overhead from these extra operations is negligible, and Cog Attention is as efficient as softmax attention. In Appendix A, we provide a detailed comparison of speed and memory usage between Cog Attention and softmax attention.
@@ 3 Enhanced Expressiveness of Cog Attention: A Mechanistic Interpretability Perspective
In this section, we provide mechanistic interpretability evidence to demonstrate that negative attention weights can enhance the expressiveness of neural networks.
3.1 Enhanced Flexibility of Attention Heads
Due to unconstrained attention weights, Cog Attention enables processes such as deletion or copying, shifting from using a static OV matrix to dynamic query-key products. This capability is a key advantage of Cog Attention, as it facilitates concurrent processes within a single head, thereby enhancing the model’s flexibility and expressiveness.
We trained a Transformer language model with 141 million parameters and a Transformer-like language model using Cog Attention of the same size, respectively. Details regarding the model training can be found in Section 4. We studied the working mechanisms of attention heads on the indirect object identification (IOI) task (Wang et al., 2023), where the model is provided with a context that includes the names of two people. For instance, given the input “Christopher and David had a lot of fun at school. David gave a ring to,” “Christopher” is the indirect object ($IO$), while “David” is the subject ($S$). The correct answer in the IOI task is always the $IO$, which in this case is “Christopher.” There are 100 samples in the dataset.
To identify the most influential attention heads contributing to correct predictions in each model, we employed the path patching algorithm (Wang et al., 2023). We identified two significant heads: the 4th Cog Attention head in Layer 9 ($\text{CH}_{9.4}$) and the 11th softmax head in Layer 9 ($\text{SH}_{9.11}$). These heads accomplish the IOI task through a process of elimination. Figure 1(a) and (b) illustrate the attention weight patterns for these heads. Additionally, we computed the attention weights from the final token to both the $IO$ and $S$ tokens, plotted against the inner product of the heads’ outputs and the individual embeddings of $IO$ and $S$, as shown in Figure 1(c) and (d).
@@ 3.1 Enhanced Flexibility of Attention Heads
In contrast, $\text{CH}_{9.4}$ assigns negative weights to $S$, effectively eliminating it while assigning minimal weights to $IO$s. Even if an $IO$ is assigned with a negative weight, we contend that the OV matrix in $\text{CH}_{9.4}$ is relieved from deletion to perform post-processing such as refinement or modification, and thus retains $IO$s. Evidence for this claim comes from the eigenvalue positivity [^3] Defined in the “Summarizing OV/QK Matrices” section in  (Elhage et al., 2021) for the OV matrices in two heads. The eigenvalue positivity, which indicates the OV matrix’s tendency toward copying (close to 1), deletion (close to -1), or abstract post-processing such as refinement or modification (close to 0), is 0.78 for $\text{CH}_{9.4}$ and -0.95 for $\text{SH}_{9.11}$, with the absolute value of $\text{CH}_{9.4}$’s eigenvalue positivity being smaller, indicating fewer contextual operations. Additionally, the low Pearson correlation coefficient of $0.13$ between the attention weights and the inner product suggests little to no correlation between these two variables; that is, $IO$s are hardly eliminated by $\text{CH}_{9.4}$.
@@
Figure 5: The output representation difference are measured for Transformer language models utilizing Cog Attention (subfigures (a) and (b)) and softmax attention subfigures (c) and (d)), respectively. Cog Attention enhances the robustness of language models against representational collapse.
3.2 More Robustness to Representational Collapse
Transformer models suffer from the issue of representational collapse. Initially, Liu et al. (2020) and Xie et al. (2023) described representational collapse as the high similarity of representations between consecutive layers. They attributed this issue to normalization layers and optimizers. Later, Barbero et al. (2024) broadened the definition, identifying another form of representational collapse that occurs within the same layer, where the representations of tokens at different positions become increasingly similar as model depth increases. They explain this by suggesting that earlier tokens have more information pathways to the final position than later tokens, leading to the “over-squashing” of information from earlier tokens into the final representation. This over-squashing causes representational collapse, making it difficult for Transformer models to distinguish between contexts that differ only slightly. Since the latter form of representational collapse is attributed to the attention mechanism—central to the focus of this paper—we adopt the definition of representational collapse as presented in (Barbero et al., 2024).
To evaluate representational collapse in Transformer-based language models, we employ two tasks, as shown in Figure 4. Task 1 “Finding a Zero” involves processing two input sequences. The first sequence consists of $n+1$ ones, while the second begins with a zero followed by $n$ ones, where $n$ varies. Given an $n$, let $\mathbf{y}^{n}_{1}$ and $\mathbf{y}^{n}_{2}$ represent the output vectors at the final position of the last layer for the first and second inputs, respectively. A higher value of $L_{+\infty}$-norm of their difference indicates that the model is better at distinguishing between the two similar contexts, reflecting reduced representational collapse (Barbero et al., 2024). In this paper, we report the relative $L_{+\infty}$-norm, defined as:
$\Delta L_{+\infty}=\frac{\lVert\mathbf{y}^{n}_{2}-\mathbf{y}^{n}_{1}\rVert_{+\infty}}{\lVert\mathbf{y}^{200}_{2}-\mathbf{y}^{200}_{1}\rVert_{+\infty}},$
because norms are not directly comparable across different models. Task 2, “Counting Ones,” employs the same evaluation approach.
We evaluate our models on these two tasks. As shown in Figure 5, for each $n$ and across both tasks, models employing Cog Attention demonstrate greater robustness against representational collapse. This advantages may be beneficial for tasks involving long and complex contexts, such as retrieval-augmented generation (RAG; Lewis et al., 2020).
Remark. This enhanced robustness against representational collapse is attributed to the negative weights in Cog Attention  which reduce redundant information pathways from earlier tokens to later positions, thereby mitigating over-squashing. A formal discussion is provided in Appendix B.